# Extremal Process of Last Progeny Modified Branching Random Walks

### v1

@@ Abstract
MSC2020 subject classification: Primary 60G55, 60G70, 60J80; Secondary 60G42, 60G50.
Key Words: Branching random walk; point process; extremal process; additive martingales; derivative martingales; regularly varying functions.
1 Introduction
Branching random walk (or BRW for short) is a particle system on the real line constructed as follows. It starts from a single particle at position 00 forming the initial generation 00 of the process. Each particle reproduces independently of all others, by creating an identically distributed point process of children around its position. We denote by $\mathcal{U}$ the set of particles in the branching random walk. For all $u\in\mathcal{U}$, we write $S_{u}$ for the position of particle $u$, $|u|$ for the generation to which it belongs and $u_{k}$ for the ancestor of $u$ alive at generation $k\leq|u|$. We only consider in the present article supercritical branching random walks, satisfying the assumption
$\operatorname{\mathbf{E}}\left(\#\{u\in\mathcal{U}:|u|=1\}\right)>1.$ (1.1)
Note that we do not make any assumption on the finiteness of #$\#\{|u|=1\}$.
For all $\theta>0$, we denote by $\kappa(\theta)=\log\operatorname{\mathbf{E}}\left(\sum_{|u|=1}e^{\theta S_{u}}\right)\in(-\infty,\infty]$ the log-Laplace transform of (the intensity measure of the reproduction law of) the BRW. Assuming that $\kappa(\theta)<\infty$, we write
$\kappa^{\prime}(\theta)=\operatorname{\mathbf{E}}\left(\sum_{|u|=1}S_{u}e^{\theta S_{u}-\kappa(\theta)}\right)$
@@ 1 Introduction
$\kappa^{\prime\prime}(\theta)=\operatorname{\mathbf{E}}\left(\sum_{|u|=1}(S_{u}-\kappa^{\prime}(\theta))^{2}e^{\theta S_{u}-\kappa(\theta)}\right)\in[0,\infty].$
Observe that by Lebesgue’s dominated convergence theorem, it is straightforward to verify that $\kappa^{\prime}(\theta)$ and $\kappa^{\prime\prime}(\theta)$ indeed correspond to the first and second derivative of $\kappa$ whenever this function is finite in a neighbourhood of $\theta$. We also remark that $\kappa^{\prime\prime}(\theta)=0$ implies that almost surely, all particles in the first generation have the same position. We bar this degenerate situation from consideration, by always assuming in this article that
$\forall x\in\mathbb{R},\quad\operatorname{\mathbf{P}}(\exists|u|=1:S_{u}\neq x)>0.$ (1.2)
Branching random walks have been the subject of a large and still expanding literature. One of the most studied features of this model is the asymptotic behaviour of the position of particles at the right tip of the BRW. Biggins [7] proved that the maximal displacement $M_{n}:=\max_{|u|=n}S_{u}$ satisfies
[…]
as long as there exists $\theta>0$ such that $\kappa(\theta)<\infty$. Addario-Berry and Reed [1] (see also Bramson and Zeitouni [10]) showed that if there exists $\theta_{0}>0$ such that
@@
and additional integrability conditions hold, then setting $m_{n}:=nv-\frac{3}{2\theta_{0}}\log n$, the sequence $(M_{n}-m_{n})$ is tight, although Hu and Shi [12] proved that this sequence exhibits almost sure fluctuations on a logarithmic scale. Aïdékon [2] proved the convergence in distribution of the centred maximal displacement under close to optimal integrability conditions for a BRW satisfying (1.3).
@@
where $\mathcal{X}_{n}=\sum_{|u|=n}\delta_{S_{u}}$ is the empirical measure of the BRW and $\tau_{x}$ is the operator corresponding to a shift by $x$ of the point measure. Madaule [16] proved that under mild conditions, the extremal process $\mathcal{Z}_{n}$ of the BRW converges in law for the topology of vague convergence of point measure to a limiting decorated Cox process $\mathcal{Z}_{\infty}$. We refer to [19] for an overview of branching random walks, and [17, 20] for background on decorated Cox processes and their connections with branching particle systems.
We take interest in this article to the last progeny modified BRW, introduced by Bandyopadhyay and Ghosh in [6], that can be constructed as follows: Let $\nu$ be a probability measure on $\mathbb{R}$ and $(Y_{u},u\in\mathcal{U})$ be a collection of i.i.d. random variables of law $\nu$, which are independent of the BRW. This model is the family of point measures defined for $n\geq 0$ by
@@
In other words, in the process $\mathcal{E}_{n}$, the position of particles at the last step $n$ in the BRW are modified by the i.i.d. random variables $(Y_{u},|u|=n)$.
The asymptotic behaviour of $\mathcal{E}_{n}$ can be described using the classical additive martingales of the branching random walk. Let $\theta>0$ such that $\kappa(\theta)<\infty$, the process defined for $n\geq 0$ by
$W_{n}(\theta)=\sum_{|u|=n}e^{\theta S_{u}-n\kappa(\theta)}$
is a non-negative martingale called the additive martingale of the BRW. Assuming that $\kappa^{\prime}(\theta)$ is finite, Biggins [8] obtained a necessary and sufficient condition for the non-degeneracy of its almost sure limit $W_{\infty}(\theta)$ (with an alternative proof by Lyons [14] based on a spine decomposition argument). More precisely, the martingale $(W_{n}(\theta))$ is uniformly integrable if and only if
$\theta\kappa^{\prime}(\theta)-\kappa(\theta)<0\text{ and }\operatorname{\mathbf{E}}(W_{1}(\theta)\log_{+}W_{1}(\theta))<\infty,$ (1.4)
where $\log_{+}(x)=\log\max(x,1)$. If (1.4) does not hold, then $W_{\infty}(\theta)=0$ a.s. Alsmeyer and Iksanov [5] obtained a more general necessary and sufficient condition for the non-degeneracy of $W_{\infty}(\theta)$ that does not depend on the well-definition of $\kappa^{\prime}(\theta)$.
Let us observe that, by assumption (1.2), the function $\theta\mapsto\kappa(\theta)$ is a strictly convex function, therefore $\theta\mapsto\theta\kappa^{\prime}(\theta)-\kappa(\theta)$ is increasing on its interval of definition. Consequently (1.4) implies that $\theta<\theta_{0}$ whenever this quantity is well-defined.
Assume that the constant $\theta_{0}$ defined in (1.3) exists. In this case, as mentioned above, the additive martingale $W_{n}(\theta_{0})$ converges to 00 a.s. However, the so-called derivative martingale defined for $n\geq 0$ by
$Z_{n}:=\sum_{|u|=n}(n\kappa^{\prime}(\theta_{0})-S_{u})e^{\theta_{0}S_{u}-n\kappa(\theta_{0})}$
plays an important role for the study of extremes of the BRW. Assuming that $\kappa^{\prime\prime}(\theta)<\infty$, Aïdékon [2] obtained necessary condition that Chen [11] proved to be sufficient for the convergence of the (signed, non-uniformly integrable) martingale $(Z_{n})$ to an a.s. positive limit $Z_{\infty}$. This necessary and sufficient conditions is
$\operatorname{\mathbf{E}}\left(W_{1}(\theta_{0})(\log_{+}W_{1}(\theta_{0}))^{2}\right)+\operatorname{\mathbf{E}}(\overline{W}_{1}\log_{+}(\overline{W}_{1}))<\infty,$ (1.5)
where $\overline{W}_{1}=\sum_{|u|=1}(\kappa^{\prime}(\theta_{0})-S_{u})_{+}e^{\theta_{0}S_{u}-\kappa(\theta_{0})}$ and $x_{+}=\max(x,0)$.
Under some integrability assumptions on the reproduction law of the BRW, Bandyopadhyay and Ghosh considered in [6] a last progeny modified BRW with perturbation law $\nu$ given by the Gumbel law of parameter $1/\theta$. Assuming that $\theta\kappa^{\prime}(\theta)-\kappa(\theta)\leq 0$, they showed that
@@
with $\mathcal{Y}_{\theta}$ a random point measure. On the one hand, if $\theta\kappa^{\prime}(\theta)-\kappa(\theta)<0$, then $m^{(\theta)}_{n}=n\frac{\kappa(\theta)}{\theta}$ and the limiting point measure $\mathcal{Y}_{\theta}$ is a Cox process (i.e. a Poisson point process with a random intensity measure) with intensity $\theta W_{\infty}(\theta)e^{-\theta x}\,\mathrm{d}x$. On the other hand, if $\theta\kappa^{\prime}(\theta)-\kappa(\theta)=0$, (1.6) holds with $m^{(\theta)}_{n}=n\kappa^{\prime}(\theta)-\frac{1}{2\theta}\log n$ and $\mathcal{Y}_{\theta}$ a Cox process with intensity $\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}\theta Z_{\infty}(\theta)e^{-\theta x}\,\mathrm{d}x$.
In [13], Kowalski obtained refined estimates on the position of the largest atom in the last progeny modified BRW for a more general class of distributions for $\nu$. The objective of the present paper is to recover the results of Bandyopadhyay and Ghosh under close to minimal conditions for the reproduction law of the BRW or for the law $\nu$. More precisely, we prove that the convergence (1.6) holds in a variety of cases for a well-chosen centring sequence $(m_{n}^{\theta})$.
@@
We endow the set of point measures on $\mathbb{R}$ with the topology of vague convergence. For $\mu$ a Radon measure on $\mathbb{R}$ and $Q$ a non-negative random variable, we denote by PPP($Q\mu$) a Cox process with intensity $Q\mu$ on $\mathbb{R}$, i.e. conditionally on $Q$ a Poisson point process with intensity $Q\mu$.
We first consider a BRW satisfying the assumptions of [5, Theorem 1.3], and prove the following result.
@@
$\nu([x,\infty))\sim Le^{-\theta x}\quad\text{as $x\to\infty$.}$ (1.7)
Then, writing $m_{n}=\frac{1}{\theta}\left(n\kappa(\theta)+\log L\right),$ the extremal process $\tau_{-m_{n}}\mathcal{E}_{n}$ converges in law to a PPP$(\theta W_{\infty}(\theta)e^{-\theta x}\,\mathrm{d}x)$.
Remark 1.2.
For the conclusion of Theorem 1.1 to hold, it is not necessary to assume that $\kappa$ is finite at any point besides $\theta$, or that $\kappa^{\prime}(\theta)>-\infty$. Integral conditions in [5, Theorem 1.3] are a necessary and sufficient condition for the convergence in distribution of $\tau_{-m_{n}}\mathcal{E}_{n}$ to hold, assuming that (1.7) is satisfied.
Adding some extra integrability conditions on the reproduction law of the BRW, we are also able to describe the asymptotic behaviour of the last progeny modified BRW for any distribution $\nu$ satisfying
$\nu([x,\infty))\sim L(x)e^{-\theta x}\quad\text{as $x\to\infty$,}$ (1.8)
where $L$ is a positive regularly varying function at $\infty$. Let us recall that a function $L$ is called regularly varying at $\infty$ with parameter $\alpha\in\mathbb{R}$ if for all $\lambda>0$, we have
$\lim_{x\to\infty}\frac{L(\lambda x)}{L(x)}=\lambda^{\alpha}.$
A function is called regularly varying at $\infty$ if there exists such a parameter $\alpha$. We refer to the book of Bingham, Goldie and Teugels [9] for background on regularly varying functions.
@@
Let $\theta>0$ such that $\kappa(\theta)<\infty$. We assume that (1.4) holds, and that there exists $\delta>0$ such that $\kappa(\theta+\delta)+\kappa(\theta-\delta)<\infty$. Let $\nu$ be a probability distribution on $\mathbb{R}$ such that there exists a regularly varying function $L$ at $\infty$ with index $\alpha$ satisfying (1.8). Then, writing
@@
the extremal process $\tau_{-m_{n}}\mathcal{E}_{n}$ converges in law to a PPP$(c_{1}\theta W_{\infty}(\theta)e^{-\theta x}\,\mathrm{d}x)$.
We now consider a situation in which $\nu$ satisfies (1.8) with the parameter $\theta_{0}$. In this situation, a similar result can be obtained, assuming a stronger condition on the function $L$ and using a modified centring term for the extremal process.

### v2

@@ Abstract.
Key words and phrases:
Branching random walk; point process; extremal process; additive martingales; derivative martingales; regularly varying functions
2010 Mathematics Subject Classification:
Primary: 60G55, 60G70, 60J80; Secondary: 60G42, 60G50.
1. Introduction
Branching random walk (or BRW for short) is a particle system on the real line constructed as follows. It starts from a single particle at position 0 forming the initial generation 0 of the process. Each particle reproduces independently of all others by creating an identically distributed point process of children around its position. We denote by $\mathcal{U}$ the set of particles in the branching random walk. For all $u\in\mathcal{U}$, we write $S_{u}$ for the position of particle $u$, $|u|$ for the generation to which it belongs and $u_{k}$ for the ancestor of $u$ alive at generation $k\leq|u|$. We only consider in the present article supercritical branching random walks, satisfying the assumption
$\operatorname{\mathbf{E}}\left[\#\{u\in\mathcal{U}:|u|=1\}\right]>1.$ (1.1)
Under this condition, the BRW survives forever with positive probability, or in other words
$\operatorname{\mathbf{P}}(\#\{u\in\mathcal{U}:|u|=n\}>0\quad\forall\,n\geq 0)>0.$
Note that we do not make any assumption on the finiteness of $\{u\in\mathcal{U}:|u|=1\}$, the random variable corresponding to the number of children of a given individual.
For all $\theta>0$, we denote by $\kappa(\theta)=\log\operatorname{\mathbf{E}}\left[\sum_{|u|=1}\mathrm{e}^{\theta S_{u}}\right]\in(-\infty,\infty]$ the log-Laplace transform of (the intensity measure of the reproduction law of) the BRW. Assuming that $\kappa(\theta)<\infty$, we write
$\kappa^{\prime}(\theta)=\operatorname{\mathbf{E}}\left[\sum_{|u|=1}S_{u}\mathrm{e}^{\theta S_{u}-\kappa(\theta)}\right]$
@@ 1. Introduction
$\kappa^{\prime\prime}(\theta)=\operatorname{\mathbf{E}}\left[\sum_{|u|=1}(S_{u}-\kappa^{\prime}(\theta))^{2}\mathrm{e}^{\theta S_{u}-\kappa(\theta)}\right]\in[0,\infty].$
By Hölder’s inequality, one immediately remarks that $\kappa$ is convex, therefore $\{\theta\in\mathbb{R}:\kappa(\theta)<\infty\}$ forms an interval. Moreover, by Lebesgue’s dominated convergence theorem, it is straightforward to verify that $\kappa^{\prime}(\theta)$ and $\kappa^{\prime\prime}(\theta)$ indeed correspond to the first and second derivative of $\kappa$ for $\theta$ in the interior of that interval. We also remark that $\kappa^{\prime\prime}(\theta)=0$ implies that almost surely, all particles in the first generation have the same position. We bar this degenerate situation from consideration by always assuming throughout this article that
$\forall x\in\mathbb{R},\quad\operatorname{\mathbf{P}}(\exists u\in\mathcal{U}:|u|=1,S_{u}\neq x)>0.$ (1.2)
Branching random walks have been the subject of a large and still expanding literature. One of the most studied features of this model is the asymptotic behaviour of the position of particles at the right tip of the BRW. Biggins (1976) proved that the maximal displacement $M_{n}:=\max_{|u|=n}S_{u}$ satisfies
$\lim_{n\to\infty}\frac{M_{n}}{n}=v:=\inf_{\theta>0}\frac{\kappa(\theta)}{\theta}\quad\text{a.s. on the survival event of the BRW,}$
as long as there exists $\theta>0$ such that $\kappa(\theta)<\infty$. Addario-Berry and Reed (2009) (see also Bramson and Zeitouni, 2009) showed that if there exists $\theta_{0}>0$ such that
@@
and additional integrability conditions hold, then setting $m_{n}:=nv-\frac{3}{2\theta_{0}}\log n$, the sequence $(M_{n}-m_{n})$ is tight, although Hu and Shi (2009) proved that this sequence exhibits almost sure fluctuations on a logarithmic scale. Aïdékon (2013) proved the convergence in distribution of the centred maximal displacement under close to optimal integrability conditions for a BRW satisfying (1.3).
@@
where $\mathcal{X}_{n}=\sum_{|u|=n}\delta_{S_{u}}$ is the empirical measure of the BRW and $\tau_{x}$ is the operator corresponding to a shift by $x$ of the point measure. Madaule (2017) proved that under mild conditions, the extremal process $\mathcal{Z}_{n}$ of the BRW converges in law for the topology of vague convergence of point measure to a limiting decorated Cox process $\mathcal{Z}_{\infty}$. We refer to Shi (2015) for an overview of branching random walks, and to Maillard and Mallein (2021) and Subag and Zeitouni (2015) for background on decorated Cox processes and their connections with branching particle systems.
In this article, we are interested in the last progeny modified BRW, introduced by Bandyopadhyay and Ghosh (2025), which can be constructed as follows: Let $\nu$ be a probability measure on $\mathbb{R}$ and $(Y_{u},u\in\mathcal{U})$ be a collection of i.i.d. random variables of law $\nu$, which are independent of the BRW. This model is the family of point measures defined for $n\geq 0$ by
@@
In other words, in the process $\mathcal{E}_{n}$, the positions of particles at the last step $n$ in the BRW are modified by the i.i.d. random variables $(Y_{u},|u|=n)$. This can be seen as a noisy observation of the positions of particles in the branching random walk.
Bandyopadhyay and Ghosh (2025) considered a last progeny modified BRW with perturbation law $\nu$ given by the Gumbel law of parameter $1/\theta$. Assuming that $\theta\kappa^{\prime}(\theta)-\kappa(\theta)\leq 0$, they showed that
@@
with $\mathcal{Y}_{\theta}$ a random point measure. On the one hand, if $\theta\kappa^{\prime}(\theta)-\kappa(\theta)<0$, then $m_{n}(\theta)=n\frac{\kappa(\theta)}{\theta}$ and the limiting point measure $\mathcal{Y}_{\theta}$ is a Cox process (i.e. a Poisson point process with a random intensity measure) with intensity $\theta W_{\infty}(\theta)\mathrm{e}^{-\theta x}\,\mathrm{d}x$. On the other hand, if $\theta\kappa^{\prime}(\theta)-\kappa(\theta)=0$, (1.4) holds with $m_{n}(\theta)=n\kappa^{\prime}(\theta)-\frac{1}{2\theta}\log n$ and $\mathcal{Y}_{\theta}$ a Cox process with intensity $\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}\theta Z_{\infty}\mathrm{e}^{-\theta x}\,\mathrm{d}x$. Here, $W_{\infty}(\theta)$ and $Z_{\infty}$ are the almost sure limits of the additive martingale $(W_{n}(\theta),n\geq 0)$ and the derivative martingale $(Z_{n},n\geq 0)$, as defined in (1.6) and (1.9), respectively. These results were extended by Kowalski (2024), who studied the asymptotic behaviour of the position of the right-most atom in the last progeny modified BRW for $\nu$ belonging to a larger class of probability distributions.
The main objective of the present article is to study the asymptotic behaviour of the full extremal process $\mathcal{E}_{n}$ as $n\to\infty$, assuming that the tail of the law $\nu$ is exponential, more precisely that there exists $\theta>0$ and a regularly varying function $L$ such that
$\nu([x,\infty))\sim L(x)\mathrm{e}^{-\theta x}\quad\text{ as $x\to\infty$.}$ (1.5)
Let us recall that a function $L$ is called regularly varying at $\infty$ with parameter $\alpha\in\mathbb{R}$ if for all $\lambda>0$, we have
$\lim_{x\to\infty}\frac{L(\lambda x)}{L(x)}=\lambda^{\alpha}.$
A function is called regularly varying at $\infty$ if there exists such a parameter $\alpha$. We refer to the book by Bingham et al. (1987) for background on regularly varying functions.
@@
We endow the set of point measures on $\mathbb{R}$ with the topology of vague convergence. For $\mu$ a Radon measure on $\mathbb{R}$ and $Q$ a non-negative random variable, we denote by PPP($Q\mu$) a Cox process with intensity $Q\mu$ on $\mathbb{R}$, i.e., conditionally on $Q$, a Poisson point process with intensity $Q\mu$.
We show in this article that the asymptotic behaviour of $\mathcal{E}_{n}$ is mainly predicted by the parameter $\theta$ in (1.5). More precisely, it satisfies a different asymptotic depending on whether $\theta<\theta_{0}$, $\theta=\theta_{0}$ or $\theta>\theta_{0}$ with $\theta_{0}$ the critical parameter of the BRW defined in (1.3). We refer to these three asymptotic regimes as subcritical, critical and supercritical, respectively, and we present our main results in each of these three regimes in the next three sections. We end the introduction with an heuristic explanation for our results.
Remark 1.1.
Aiming for generality of our results, we do not wish to work under the assumption that the parameter $\theta_{0}$ exists except if necessary. However, by strict convexity of $\kappa$, we remark that $\theta\mapsto\theta\kappa^{\prime}(\theta)-\kappa(\theta)$ is increasing whenever it is well-defined. Therefore, we will determine if the last progeny modified BRW is in the subcritical, critical or supercritical regime by observing the sign of $\theta\kappa^{\prime}(\theta)-\kappa(\theta)$.
1.1. Extremal process in the subcritical regime
Let $\nu$ be a probability distribution satisfying (1.5) and $S$ a BRW such that $\kappa(\theta)<\infty$. In the subcritical regime, we claim that the asymptotic behaviour of $\mathcal{E}_{n}$ can be described using the so-called additive martingale of the BRW, defined as follows:
$W_{n}(\theta):=\sum_{|u|=n}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}.$ (1.6)
It is immediate from its definition that $(W_{n}(\theta),n\geq 0)$ is a non-negative martingale, therefore converges almost surely to a limit $W_{\infty}(\theta)$. Assuming that $\kappa^{\prime}(\theta)$ is well-defined, Biggins (1977) obtained a necessary and sufficient condition for its uniform integrability (with an alternative proof by Lyons, 1997 based on a spine decomposition argument). More precisely, if $\kappa^{\prime}(\theta)$ is well-defined, then the martingale $(W_{n}(\theta),n\geq 0)$ is uniformly integrable if and only if
$\theta\kappa^{\prime}(\theta)-\kappa(\theta)<0\text{ and }\operatorname{\mathbf{E}}\left[W_{1}(\theta)\log_{+}W_{1}(\theta)\right]<\infty,$ (1.7)
where $\log_{+}(x)=\log\max(x,1)$. If (1.7) does not hold, then $W_{\infty}(\theta)=0$ a.s. This result was extended by Alsmeyer and Iksanov (2009), who obtained a necessary and sufficient condition for the uniform integrability of $(W_{n}(\theta),n\geq 0)$ regardless of the well-definition of $\kappa^{\prime}(\theta)$.
For the last progeny modified BRW, we show that the asymptotic behaviour of its extremal process is in the subcritical regime whenever $(W_{n}(\theta),n\geq 0)$ is uniformly integrable, i.e., as soon as the conditions of Alsmeyer and Iksanov are satisfied. However, the more restrictive condition (1.7) allows us to work with a larger class of perturbation distribution $\nu$. We therefore state two versions of our results in the subcritical regime, the first one holding under minimal conditions on the BRW but with stronger assumptions on $\nu$, the second under minimal conditions for the law $\nu$ but with the more restrictive conditions on the BRW.
We first consider a BRW satisfying the assumptions of (Alsmeyer and Iksanov, 2009, Theorem 1.3) and a perturbation distribution $\nu$ having an exact exponential tail.
@@ 1.1. Extremal process in the subcritical regime
$\nu([x,\infty))\sim L\mathrm{e}^{-\theta x}\quad\text{as $x\to\infty$.}$ (1.8)
Then, writing
$m_{n}=n\frac{\kappa(\theta)}{\theta}+\frac{1}{\theta}\log L,$
[…]
Next, we consider a BRW satisfying (1.7) and a perturbation distribution $\nu$ satisfying (1.5).
@@
Let $\theta>0$ such that $\kappa(\theta)<\infty$. We assume that there exists $\delta>0$ such that $\kappa(\theta+\delta)+\kappa(\theta-\delta)<\infty$ and that (1.7) holds. Let $\nu$ be a probability distribution on $\mathbb{R}$ satisfying (1.5) with a regularly varying function $L$ of index $\alpha$. Then, writing
@@
[…]
Noting that a constant is a regularly varying function of index 0, we see that these two statements are consistent. Theorem 1.2 corresponds to the case $\alpha=0$ and $c_{1}=1$; however, for its conclusion to hold, it is not necessary to assume that $\kappa$ is finite at any point other than $\theta$, or that $\kappa^{\prime}(\theta)>$-$\infty$, unlike in Theorem 1.3. The conditions we worked under in these results were chosen to obtain an explicit centring sequence $(m_{n})$. However, we believe that it is a generic result in the subcritical regime that there will exist a centring sequence such that $\tau_{-m_{n}}\mathcal{E}_{n}$ converges in law to a Poisson point process with intensity $W_{\infty}(\theta)\mathrm{e}^{-\theta x}\mathrm{d}x$.
1.2. Extremal process in the critical regime
Let $S$ be a BRW, we assume that there exists $\theta_{0}>0$ such that (1.3) holds, and that $\nu$ satisfies (1.5) with $\theta=\theta_{0}$. We refer to this situation as the critical regime. We observe that the martingale $(W_{n}(\theta_{0}),n\geq 0)$ converges to 0 almost surely, by the previously mentioned result of Biggins. However, the so-called derivative martingale defined for $n\geq 0$ by
$Z_{n}:=\sum_{|u|=n}(n\kappa^{\prime}(\theta_{0})-S_{u})\mathrm{e}^{\theta_{0}S_{u}-n\kappa(\theta_{0})}$ (1.9)
works as a good replacement for this martingale. Immediate computations show that $(Z_{n},n\geq 0)$ is a signed martingale, which is generally not uniformly integrable. Assuming that $\kappa^{\prime\prime}(\theta)<\infty$, Aïdékon (2013) obtained necessary condition that Chen (2015) proved to be sufficient for the convergence of $Z_{n}$ to an a.s. non-negative non-degenerate limit $Z_{\infty}$. This necessary and sufficient condition is
$\operatorname{\mathbf{E}}\left[W_{1}(\theta_{0})(\log_{+}W_{1}(\theta_{0}))^{2}\right]+\operatorname{\mathbf{E}}\left[\overline{W}_{1}\log_{+}(\overline{W}_{1})\right]<\infty,$ (1.10)
where $\overline{W}_{1}=\sum_{|u|=1}(\kappa^{\prime}(\theta_{0})-S_{u})_{+}\mathrm{e}^{\theta_{0}S_{u}-\kappa(\theta_{0})}$ and $x_{+}=\max(x,0)$.
Under these conditions, as well as a slightly more stringent condition on the tail of $\nu$, we obtain a similar result to the subcritical regime, with $Z_{\infty}$ playing the role of $W_{\infty}(\theta)$, and a modified centring term for the extremal process.